OpenStudy (anonymous):

Solve the formula P = 17.00 + 0.75h for h. Explain your steps for solving the formula for h.

OpenStudy (istim):

There's no P here now.

OpenStudy (istim):

It's a little difficult to solve with 2 variables.

OpenStudy (anonymous):

0.75h = P - 17.00 h = (p-17)/0.75

OpenStudy (anonymous):

You solve for h in terms of P

OpenStudy (istim):

Oh. Makes sense. So Hades got it then.